# Should You Use Your Large Language Model to Explore or Exploit?

### v2

@@ A.3 Additional details for text-based CB puzzles
  • The expected reward for petting the animal is 0.01 if it is a bear, 0.7 if it is a dog, 0.3 if it is a cat and the time of day is morning or afternoon, 0.7 if it is a cat and the time of day is evening or night, and otherwise 0.5.
  • The expected reward for leaving the room is always 0.5
  • If the animal is a bear, the expected reward for using the tool is 0.1. Otherwise, if the tool is a key and the table item is a chest, the expected reward is 0.9. Otherwise, it is 0.4
  • If the animal is a bear, the expected reward for eating food is always 0.5. Otherwise, the expected reward is 0.8 for cake, 0.6 for an apple, 0.2 for a nut, and 0.5 otherwise.
  • If the animal is a bear, the expected reward for pressing the button is 0.1. Otherwise if the button is green and the time of day is morning, or the button is yellow and the time of day is afternoon, or the button is orange and the time of day is evening, or the button is red and the time of day is night, then the expected reward is 0.9. In all other cases, the expected reward is 0.25.
A.4 Robustness Check with Other Models
@@
Appendix B Appendix for Section 3: LLMs as exploration oracles
B.1 Prompts
“All-at-once” prompt with encouragement for open-ended question puzzle:
@@ B.1 Prompts
“One-by-one” prompt without encouragement for arXiv puzzle:
@@ A Kernel Score Perspective on Forecast Disagreement and the Linear Pool
B.3 Additional results for explore puzzles
@@
Below are the individual plots for the remaining $9$ questions using the Sentence-BERT encoder.
Figure 24: Results on Question 1
Figure 25: Results on Question 2
Figure 26: Results on Question 3
Figure 27: Results on Question 4
Figure 28: Results on Question 5
Figure 29: Results on Question 6
Figure 30: Results on Question 7
Figure 31: Results on Question 8
Figure 32: Results on Question 9
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.39 0.4 0.4 0.4 0.41 0.42 0.56
Q1: 0.63 0.66 0.66 0.65 0.63 0.64 0.57
Q2: 0.41 0.43 0.41 0.4 0.36 0.41 0.41
Q3: 0.5 0.49 0.6 0.67 0.67 0.66 0.78
Q4: 0.33 0.46 0.4 0.43 0.38 0.42 0.42
Q5: 0.39 0.41 0.47 0.3 0.31 0.42 0.43
Q6: 0.45 0.78 0.78 0.79 0.78 0.78 0.75
Q7: 0.84 0.84 0.84 0.84 0.84 0.83 0.83
Q8: 0.39 0.41 0.47 0.36 0.37 0.37 0.39
Q9: 0.47 0.46 0.43 0.46 0.48 0.45 0.44
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.4 0.4 0.34 0.35 0.37 0.52 0.55
Q1: 0.45 0.51 0.66 0.61 0.58 0.56 0.59
Q2: 0.36 0.4 0.41 0.4 0.4 0.37 0.43
Q3: 0.48 0.42 0.41 0.49 0.4 0.6 0.49
Q4: 0.5 0.52 0.52 0.51 0.5 0.51 0.49
Q5: 0.4 0.39 0.39 0.43 0.45 0.4 0.41
Q6: 0.74 0.67 0.79 0.79 0.77 0.55 0.64
Q7: 0.69 0.82 0.78 0.79 0.79 0.77 0.8
Q8: 0.45 0.5 0.47 0.56 0.55 0.42 0.51
Q9: 0.44 0.39 0.37 0.4 0.47 0.46 0.47
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.33 0.4 0.4 0.43 0.43 0.47 0.52
Q1: 0.62 0.62 0.62 0.62 0.61 0.61 0.6
Q2: 0.43 0.43 0.44 0.43 0.43 0.44 0.45
Q3: 0.58 0.57 0.59 0.61 0.59 0.6 0.6
Q4: 0.36 0.41 0.47 0.48 0.47 0.46 0.45
Q5: 0.39 0.48 0.49 0.48 0.48 0.49 0.49
Q6: 0.83 0.81 0.82 0.81 0.81 0.8 0.79
Q7: 0.8 0.81 0.81 0.81 0.81 0.8 0.79
Q8: 0.38 0.42 0.44 0.44 0.47 0.5 0.5
Q9: 0.45 0.47 0.47 0.46 0.45 0.45 0.43
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.35 0.37 0.5 0.54 0.55 0.58 0.59
Q1: 0.55 0.58 0.6 0.61 0.61 0.61 0.62
Q2: 0.45 0.45 0.44 0.44 0.45 0.45 0.42
Q3: 0.55 0.55 0.54 0.53 0.53 0.53 0.54
Q4: 0.33 0.49 0.48 0.49 0.49 0.48 0.46
Q5: 0.4 0.43 0.44 0.44 0.45 0.45 0.45
Q6: 0.83 0.83 0.82 0.82 0.81 0.81 0.8
Q7: 0.75 0.8 0.81 0.82 0.81 0.82 0.82
Q8: 0.35 0.4 0.45 0.47 0.49 0.5 0.53
Q9: 0.49 0.49 0.5 0.49 0.5 0.47 0.48
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: -0.01 -0.0 0.01 0.01 0.03 0.03 0.02
Q1: 0.01 0.01 0.02 0.02 0.04 0.04 0.03
Q2: 0.01 -0.0 0.03 0.02 0.02 0.03 0.02
Q3: 0.05 0.04 0.01 0.05 0.04 0.03 0.04
Q4: 0.02 0.01 0.03 0.02 0.02 0.04 0.01
Q5: -0.02 -0.01 0.02 0.03 0.02 0.03 0.05
Q6: 0.0 0.0 0.03 0.03 0.04 0.02 0.03
Q7: 0.01 0.05 0.03 0.04 0.04 0.02 0.04
Q8: -0.04 0.03 0.04 0.04 0.03 0.03 0.04
Q9: 0.04 0.0 0.04 0.03 0.04 0.03 0.04
@@
Figure 33: Results for arXiv category econ.GN
Figure 34: Results for arXiv category eess.SP
Figure 35: Results for arXiv category q-bio.BM
Figure 36: Results for arXiv category math.ST
B.4 Benchmarking encoders
@@
B.5 Robustness Check with Other Models
In this appendix, we include results for Qwen2.5-7B-Instruct (Qwen-2.5), Gemma 3 12B (Gemma-3), and Mistral-7B-Instruct-v0.3 (Mistral-7B) on our explore Q/A puzzles. As was the case with the GPT models, performance increased with $K$ when the candidate answers are generated one-by-one. When generating answers all-at-once, we observed that Qwen-2.5 and Mistral-7B had trouble generating concise answers, leading to a drop in performance as $K$ increases.
Figure 37: Results for Qwen-2.5 in the setting of Figure 6.
Figure 38: Results for Mistral-7B in the setting of Figure 6.
Figure 39: Results for Gemma-3 in the setting of Figure 6.

### v3

@@ B.3 Additional details for text-based CB puzzles
  • ExpRev for petting the animal is 0.01 if it is a bear, 0.7 if it is a dog, 0.3 if it is a cat and the time of day is morning or afternoon, 0.7 if it is a cat and the time of day is evening or night, and otherwise 0.5.
  • ExpRev for leaving the room is always 0.5
  • If the animal is a bear, ExpRev for using the tool is 0.1. Otherwise, if the tool is a key and the table item is a chest, ExpRev is 0.9. Otherwise, it is 0.4
  • If the animal is a bear, ExpRev for eating food is always 0.5. Otherwise, ExpRev is 0.8 for cake, 0.6 for an apple, 0.2 for a nut, and 0.5 otherwise.
  • If the animal is a bear, ExpRev for pressing the button is 0.1. Otherwise if the button is green and the time of day is morning, or the button is yellow and the time of day is afternoon, or the button is orange and the time of day is evening, or the button is red and the time of day is night, then ExpRev is 0.9. In all other cases, ExpRev is 0.25.
B.4 Robustness Check with Other Models
@@
Figure 23: Performance of DeepSeek-R1 on our numerical CB puzzle.
Appendix C Details for Section 3: LLMs as exploration oracles
C.1 Prompts
Prompt for the MovieLens experiments:
[SYSTEM] You are a movie expert helping a user choose a movie.
[USER] Here is a list of movies with their numeric IDs: …
From this list, choose {K} movies. You don’t know what taste in movies the user has, so select a diverse set of movies from different genres such that they will most likely enjoy at least one of the movies you select. Respond ONLY with the {K} numeric IDs, one per line, with no extra text. Do NOT consider a movie’s popularity when deciding whether to select it.
For the arXiv and Q/A tasks, we used a 2x2 space of prompts, as discussed in Section 3: $\left\{\,\text{all-at-once, 1-by-1}\,\right\}\times\left\{\,\text{with, without}\,\right\}$ encouragement. We provide two illustrative examples below.
“All-at-once” prompt with encouragement for the Q/A task:
@@ C.1 Prompts
“One-by-one” prompt without encouragement for the arXiv task:
@@ A Kernel Score Perspective on Forecast Disagreement and the Linear Pool
C.4 Robustness Check: Q/A task with another embedding model
@@
C.5 Additional Results: arXiv task
Figure 25: Results for other arXiv categories
@@ C.5 Additional Results: arXiv task
C.6 Additional Results: QA task
While previously we presented the plot for Question 0 (in Figure 7), below are the individual plots for the remaining $9$ questions, in the same notation. Throughout, we used the Sentence-BERT encoder.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.39 0.4 0.4 0.4 0.41 0.42 0.56
Q1: 0.63 0.66 0.66 0.65 0.63 0.64 0.57
Q2: 0.41 0.43 0.41 0.4 0.36 0.41 0.41
Q3: 0.5 0.49 0.6 0.67 0.67 0.66 0.78
Q4: 0.33 0.46 0.4 0.43 0.38 0.42 0.42
Q5: 0.39 0.41 0.47 0.3 0.31 0.42 0.43
Q6: 0.45 0.78 0.78 0.79 0.78 0.78 0.75
Q7: 0.84 0.84 0.84 0.84 0.84 0.83 0.83
Q8: 0.39 0.41 0.47 0.36 0.37 0.37 0.39
Q9: 0.47 0.46 0.43 0.46 0.48 0.45 0.44
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.4 0.4 0.34 0.35 0.37 0.52 0.55
Q1: 0.45 0.51 0.66 0.61 0.58 0.56 0.59
Q2: 0.36 0.4 0.41 0.4 0.4 0.37 0.43
Q3: 0.48 0.42 0.41 0.49 0.4 0.6 0.49
Q4: 0.5 0.52 0.52 0.51 0.5 0.51 0.49
Q5: 0.4 0.39 0.39 0.43 0.45 0.4 0.41
Q6: 0.74 0.67 0.79 0.79 0.77 0.55 0.64
Q7: 0.69 0.82 0.78 0.79 0.79 0.77 0.8
Q8: 0.45 0.5 0.47 0.56 0.55 0.42 0.51
Q9: 0.44 0.39 0.37 0.4 0.47 0.46 0.47
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.33 0.4 0.4 0.43 0.43 0.47 0.52
Q1: 0.62 0.62 0.62 0.62 0.61 0.61 0.6
Q2: 0.43 0.43 0.44 0.43 0.43 0.44 0.45
Q3: 0.58 0.57 0.59 0.61 0.59 0.6 0.6
Q4: 0.36 0.41 0.47 0.48 0.47 0.46 0.45
Q5: 0.39 0.48 0.49 0.48 0.48 0.49 0.49
Q6: 0.83 0.81 0.82 0.81 0.81 0.8 0.79
Q7: 0.8 0.81 0.81 0.81 0.81 0.8 0.79
Q8: 0.38 0.42 0.44 0.44 0.47 0.5 0.5
Q9: 0.45 0.47 0.47 0.46 0.45 0.45 0.43
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.35 0.37 0.5 0.54 0.55 0.58 0.59
Q1: 0.55 0.58 0.6 0.61 0.61 0.61 0.62
Q2: 0.45 0.45 0.44 0.44 0.45 0.45 0.42
Q3: 0.55 0.55 0.54 0.53 0.53 0.53 0.54
Q4: 0.33 0.49 0.48 0.49 0.49 0.48 0.46
Q5: 0.4 0.43 0.44 0.44 0.45 0.45 0.45
Q6: 0.83 0.83 0.82 0.82 0.81 0.81 0.8
Q7: 0.75 0.8 0.81 0.82 0.81 0.82 0.82
Q8: 0.35 0.4 0.45 0.47 0.49 0.5 0.53
Q9: 0.49 0.49 0.5 0.49 0.5 0.47 0.48
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: -0.01 -0.0 0.01 0.01 0.03 0.03 0.02
Q1: 0.01 0.01 0.02 0.02 0.04 0.04 0.03
Q2: 0.01 -0.0 0.03 0.02 0.02 0.03 0.02
Q3: 0.05 0.04 0.01 0.05 0.04 0.03 0.04
Q4: 0.02 0.01 0.03 0.02 0.02 0.04 0.01
Q5: -0.02 -0.01 0.02 0.03 0.02 0.03 0.05
Q6: 0.0 0.0 0.03 0.03 0.04 0.02 0.03
Q7: 0.01 0.05 0.03 0.04 0.04 0.02 0.04
Q8: -0.04 0.03 0.04 0.04 0.03 0.03 0.04
Q9: 0.04 0.0 0.04 0.03 0.04 0.03 0.04
C.7 Robustness Check with Other Models
We include results for Qwen2.5-7B-Instruct (Qwen-2.5), Gemma 3 12B (Gemma-3), and Mistral-7B-Instruct-v0.3 (Mistral-7B) on our explore Q/A puzzles, all in the setting of Figure 7. As was the case with the GPT models, performance increased with $K$ when the candidate answers are generated one-by-one. When generating answers all-at-once, we observed that Qwen-2.5 and Mistral-7B had trouble generating concise answers, leading to a drop in performance as $K$ increases.
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
C.8 Benchmarking encoders
@@
C.9 Explore Experiments on the Movie Lens Dataset
Prompt:
[SYSTEM] You are a movie expert helping a user choose a movie.
[USER] Here is a list of movies with their numeric IDs:{movie_list_str}
From this list, choose K movies. You don’t know what taste in movies the user has, so select a diverse set of movies from different genres such that they will most likely enjoy at least one of the movies you select. Respond ONLY with the K numeric IDs, one per line, with no extra text. Do NOT consider a movie’s popularity when deciding whether to select it.